OpenStudy (anonymous):

!HELP! minerva wants to mail a package at he post office they charge $1.50 fot he first five ounces and $.22 per ounce for the remainder of the weight, if her package weights two pounds what change will she recieve from $10?

OpenStudy (anonymous):

2 pounds = 2 * 16 = 32 ounces 32 - 5 = 27 ounces 27 * 0.22 = 5.94 dollars 10 - (1.50 + 5.94) = ?

OpenStudy (anonymous):

14.44?

OpenStudy (anonymous):

Nah. 10 - (1.50 + 5.94) = 10 -1.5 - 5.94 So the answer is 2.56 dollars.

OpenStudy (anonymous):

yeha sorry i thought it was a plus sign but thanx tho

OpenStudy (anonymous):

a minus" sign jajaj